OpenStudy (anonymous):

Three cards are chosen at random from a deck without replacement. What is the probability of getting a jack, a ten, and a nine? (Hint: reduce each fraction separately before multiplying)

OpenStudy (anonymous):

There are 4 Jacks, 4 tens, and 4 nines in a standard deck so the probablility for drawing 1 of any of them is 4/52 or 1/13

OpenStudy (anonymous):

Now to find the combined probability you would do mulitplication

OpenStudy (anonymous):

P(10) = 4/48 P(jack) = 4/52 I think :)

OpenStudy (anonymous):

what about the 9

OpenStudy (anonymous):

You know the probability of drawing a Jack is 1/13 or 4/52, but when you draw a ten the next time, probability changes to 4/51 <---- because you didnt replace it

OpenStudy (anonymous):

And same with drawing a nine, the probability for nine is now 4/50 or 2/25

OpenStudy (anonymous):

If you multiply 1/13, 4/51/ and 2/25 that will be your answer

OpenStudy (anonymous):

thank you